OpenStudy (anonymous):

Find the area of a triangle whose base is 2√3 in. and height is 3√3 in. I dont understand square roots can you plz explain the basic part...

OpenStudy (anonymous):

\[(\sqrt{2})^2 =2\] \[(\sqrt{3})^2 =3\] Area = (1/2)x base x height

OpenStudy (anonymous):

\[\sqrt {a} \times \sqrt b = \sqrt{a \times b}\] \[\sqrt a \times \sqrt a = \sqrt{a \times a} = a\]

OpenStudy (anonymous):

i.e \[\sqrt 6 \times \sqrt 2 = \sqrt{6 \times 2} = \sqrt{12}\]

OpenStudy (anonymous):

do you get it @Lilly95 ?

OpenStudy (anonymous):

im confused how did you get 6 and 2

OpenStudy (anonymous):

those were just examples

OpenStudy (anonymous):

i.e. means something like for example

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

in your problem area = bh/2 therefore \[A = \frac{2\sqrt 3 \times 3 \sqrt 3}{2}\] \[2\sqrt 3 \times 3 \sqrt 3 = (2\times 3)(\sqrt{3 \times 3}) = 6(3) = 18\]

OpenStudy (anonymous):

you ahve to divide it by 2 so \[A = \frac{18}{2} = 9\]

OpenStudy (anonymous):

you understand right @Lilly95

OpenStudy (anonymous):

Know i do

OpenStudy (anonymous):

Thnx